SPISAK NAUČNIH RADOVA

1. Atanasijević, I., Milogradov, J.: ”Posmatranja Sunčeve korone prilikom potpunog pomračenja Sunca 1961. godine”,Vesnik Društva matematičara i fizičara SR Srbije, XV (1-4), 1963, pp. 161-169.

2. Atanasijević, I., Milogradov, J.: ”Struktura Sunčeve korone 15. februara 1961. godine”, Publikacije Astronomske opservatorije u Beogradu, 12, 1968, pp. 143- 156.

3. Milogradov-Turin, J.: ”Structure of the continuum background at high galactic latitudes at 38 MHz”, Publications of the Department of Astronomy, Faculty of Sciences, University of Beograd, 2, 1970, pp.5-9.

4. Milogradov-Turin, J.: ”The distortion of the galactic plane emission in the anticentre region”, Publications of the Department of Astronomy, Faculty of Sciences, University of Beograd, 3, 1971, pp. 25-28.

5. Milogradov-Turin, J.: ”The extension of Loop III into southern latitude”, Monthly Notices of the Royal Astronomical Society, 157, 1972, pp. 1P-4P.

6. Milogradov-Turin, J.: ”The survey of the region 6h> α >19h, −25^{∘} < δ < 70^{∘} at 38 MHz”,M.Sc. Thesis, University of Manchester, l972, pp. 1-164.

7. Milogradov-Turin, J.; Smith, F. G.: ”A survey of the radio background at 38 MHz”, Monthly Notices of the Royal Astronomical Society, 161, 1973, pp. 269-279.

8. Milogradov-Turin, J.: ”Pregled neba u oblasti 6h> α > 19h, - 25^{∘} < δ < 70^{∘} na 38 MHz”, Publikacije Astronomske opservatorije u Beogradu, 20, 1975, pp. 27-28 (rezime magistarske teze).

9. Milogradov-Turin, J.: ”The spectrum of the background radiation between 38 and 404 MHz”,Memorie della Societa Astronomica Italiana (Proceedings of the II ERAM, Trieste, 1974), 45, 1974, pp. 85-87.

10. Milogradov-Turin, J.; Ninković, S.: ”The signature of the galactic radio halo between 38 MHz and 404 MHz”, Publications de l’ observatoire astronomique de Beograd, 26, l979, pp. 133-137.

11. Milogradov-Turin, J.: ”Absorption of radio waves in HII regions S54, S49 and S45”,Publications of the Astronomical Observatory of Sarajevo, 1, 1981, pp. 233-245.

12. Milogradov-Turin, J.: ”Thermal and nonthermal components of the galactic radiation at 38 MHz for 46^{∘} < l < 65^{∘}, b = 0^{∘}”, Publications of the Department of Astronomy, Faculty of Sciences, University of Beograd, 11, 1981, pp. 27-34.

13. Milogradov-Turin, J.:”Pregled neba na 38 MHz i galakticki spektralni indeks”, Doktorska disertacija, Univerzitet u Beogradu, 1982, pp. 1-473.

14. Milogradov-Turin, J.: ”Cyg X at 38 MHz”,Hvar Observatory Bulletin Supplement, 6, 1982, pp. 23-33.

15. Milogradov-Turin, J.: ”The flux density of Cyg A at 38 MHz”,Hvar Observatory Bulletin Supplement, 6, l982, pp.35-37.

16. Milogradov-Turin, J.: ”Loop I (The North Polar Spur Region) – A Quasy Radio Halo”,Proceedings of the 106th Symposium of the I.A.U. ”The Milky Way Galaxy”, Groningen-l983, l985, pp. 245-246.

17. Milogradov-Turin, J.: ”Radio Observations of the North Equatorial Cap at 38MHz”, Monthly Notices of the Royal Astronomical Society, 208, l984, pp. 379-380.

18. Milogradov-Turin, J.: ”A survey of the Radio Background at 38 MHz and the galactic spectral index”, Publication de l’observatoire astronomique de Belgrade, 33, 1985, pp. 28-29. (rezime doktorske teze).

19. Milogradov-Turin, J.: ”The Ionospheric Refraction at 38 MHz”, Publications of the Department of Astronomy, University of Beograd, 14, 1986, pp. 33-36.

20. Milogradov-Turin, J.:”The 4V (four vectors) Model of the Radio Emission”,Proceedings of the Tenth E.R.A.M.,Prague, l987, Publications of the Astronomical Insitute of the Czechoslovak Academy of Sciences, 69, 1987, pp. 225-229.

21. Milogradov-Turin, J. ”Cosmic Synchrotron Radiation”, Proceedings of the II Workshop ”Astrophysics in Yugoslavia”, Beograd, 1987, pp. 7-8.

22. Milogradov-Turin, J. ”The possibilities of the development of the non optical astronomy in Yugoslavia”, Proceedings of the III Workshop ”Astrophysics in Yugoslavia”, Zagreb, 1989, pp.

23. Milogradov-Turin, J. ”Temperature parameter as a proof for the absence of an elliptical galactic halo”, Proceedings of the IAU Symposium No 144. ”The Interstellar Disk-Halo Connection In Galaxies, Leiden, 1990, p. 23-24.

24. Milogradov-Turin, J. and Nikolić, S.: ”Differential spectral indices of the radio loops between 38, 408 and 1420 MHz”, Bull. Astron. Belgrade, No 151, 7-11, 1995.

25. Milogradov-Turin, J. and Nikolić, S.: ” T-T Plots of Radio Spurs between 38, 408 and 1420 MHz” Publ. Obs. Astron. Belgrade, No 49, 175 -178, 1995.

26. Milogradov-Turin, J. and Nikolić, S.: ”The Quasihysteresis Effect in T- T Plots of Radio Spurs between 38, 408 and 1420 MHz” Bull. Astron. Belgrade, No 152, 5-9, 1995.

27. Milogradov-Turin, J. and Nikolić, S.: ”The Shape of Radio Spectra of Spurs between 38, 408 and 1420 MHz” Bull. Astron. Belgrade, No 152, 11- 15, 1995.

28. Milogradov-Turin, J.: ”Rasprava o meteoritima Atanasija Stojkovića”, Zbornik radova naučnog skupa ”Prirodne i matematičke nauke u Srba u 18. i u prvoj polovini 19. veka”, SANU – Univerzitet u Novom Sadu – Matica srpska, Novi Sad, 49-56, 1995.

29. Milogradov-Turin, J.: ”Bibliografija knjige ”O vazdušnim kamenovima i njihovom poreklu” Atanasija Stojkovića”, Zbornik radova naučnog skupa ”Prirodne i matematičke nauke u Srba u 18. i u prvoj polovini 19. veka”, SANU – Univerzitet u Novom Sadu – Matica srpska, Novi Sad, 57-61, 1995.

30. Milogradov-Turin, J., Urošević, D.: ”The Small Circles of the Origem and the Monoceros Loops at 1420 MHz”, Publ. Obs. Astron. Belgrade, No 54, 47-48, 1996.

31. Milogradov-Turin, J., Urošević, D.: ”Geometry of Large Radio Loops at 1420 MHz”, Bull. Astron. Belgrade, No 155, 41-45, 1997.

32. Milogradov-Turin, J.: ”Jedan prilog istoriji Katedre za astronomiju Matematičkog fakulteta u Beogradu”, Zbornik radova sa skupa ”Razvoj astronomije kod Srba”, Publ. Obs. Astron. Belgrade, No 56, 65-75, 1997.

33. Milogradov-Turin, J.: ”Astronomija na seminarima za nastavnike srednjih škola”, Zbornik radova sa skupa ”Razvoj astronomije kod Srba”, Publ. Obs. Astron. Belgrade, No 56, 229-247, 1997.

34. Milogradov-Turin, J.: ”Chair of Astronomy of the University of Belgrade”, Proceedings of the 4^{th} Yugoslav-Romanian Astronomical Meeting”, Publ. Obs. Astron. Belgrade, No 60, 193-195, 1998.

35. Urošević, D., Milogradov-Turin, J.: ”The Distances to the Origem and the Monoceros Loops ”, Serb. Astron. J., No 157, 35-38, 1998.

36. Simovljević, J., Milogradov-Turin, J.: ”Astronomija” , u ”Spomenica 125 godina Matematičkog fakulteta” (Ur. N. Bokan), Univerzitet u Beogradu Matematički fakultet, 59-105, Beograd, 1998, isto na CD.

37. Milogradov-Turin, J.: ”Hill Named Stojković”, u Proceedings of the XII National Conference of Yugoslav Astronomers and International Workshop on the Development of Astronomical Databases (Ed. L. Popović and M. Dačić), Publ. Obs. Astron. Belgrade, No 65, 173-176, 1999.

38. Milogradov-Turin, J.: ”Astronomers of Serbia in Professional Societies (II): The Union of Societies of Mathematicians and Physicists on the Occasion of 50 Years Since Its Foundation”, u Proceedings of the XII National Conference of Yugoslav Astronomers and International Workshop on the Development of Astronomical Databases (Ed. L. Popović and M. Dačić), Publ. Obs. Astron. Belgrade, No 65, 177-184, 1999.

39. Milogradov-Turin, J.: ”Prilozi za biografiju Milutina Milankovića”, u Simpozijum Milanković – juče, dans, sutra (Ur. V. Milićević), Beograd, 125-143, 1999.

40. Milogradov-Turin, J.: ”Pola veka od osnivanja Saveza društava matematičara i fizičara Jugoslavije”, Zbornik radova sa 10. kongresa fizičara Jugoslavije, Vrnjačka banja, 27-29. mart 2000 (Ur. B. Milić i D. Markušev), 1021-1030.

41. Milogradov-Turin, J.: : ”Brdo Stojković i istorija vazdušnog kamenja” , Flogiston, god. 7, br. 11, 15-46, 2001

42. Milogradov-Turin, J., Mićić: ”Fifty Years of The Union of The Societies of Mathematicians, Physicists and Astronomers of Yugoslavia – a Year After”, 2001.